OpenStudy (anonymous):

how do i solve the inequality -6<(-1/2)(-6)-4

OpenStudy (jdoe0001):

use PEMDAS :)

OpenStudy (anonymous):

I tried that i got it wrong 3 times this is my 4th time resubmitting it so i really want to get it right this time

OpenStudy (jdoe0001):

\(\bf -6 < \left(-\cfrac{1}{2}\right)(-6)(-4)\implies -6 < -\cfrac{1}{2}\times -6\times -4 \)

OpenStudy (jdoe0001):

multiplying from LEFT TO RIGHT on the right-hand side, we get \(\bf -\cfrac{1}{2}\times -6\times -4\implies -\cfrac{-6\times -4}{2}\implies -\cfrac{24}{2}\)

OpenStudy (jdoe0001):

we do not do the left side, since it's just a constant already

OpenStudy (jdoe0001):

\(\bf -\cfrac{24}{2}\implies -12\) so we end up with -6 < -12 which is not an inequaliity, because, -12 is not greater than -6 recall that for negative numbers, the closer to 0, the greater the number notice that -6 is closer to 0 than -12

OpenStudy (jdoe0001):

not a valid inequality for that matter